How many ounces of an alloy containing 30% gold on the world market must be added to 10 ounces of an alloy containing 5% to produce an alloy containing 20% gold?
Found 2 solutions by galactus, modi.kk91:
Answer by galactus(183)

Let x be the amount of the 30% being added.
So, we have
.30x+.05(10)=.20(10+x)
.30x+.5=2+.20x
.10x=1.5
x=1.5/.10=15
Add 15 ounces of the 30% alloy

Answer by modi.kk91(20)

bye method of allegation mixture= n1/n2=A2-A/A-A1

A1= 5% A2= 30%
20%

30-20 20-5
n1 =10 n2 =15

ratio =10:15=2:3
now 2:3::10:x
2/3=10/x
then
x=15